# Learning Continually by Spectral Regularization

### v1

@@ 4 Regularizing for Trainability in Continual Learning
If important properties of the initialization are lost during the course of learning, it is natural to regularize the neural network towards the initialization. This is the motivation for regularization [17] and weight reinitialization [23, 6] as loss of plasticity mitigators. However, our motivation is to more directly target key properties of initialization, using the insights from Section 3.
@@
To directly preserve key initialization properties, we will develop regularizers $\mathcal{R}_{\tau(t)}(\theta,s)$ that are a function of (i) the parameters of the neural network, $\theta$, (ii) the data, through the current task $\tau(t)$, and of (iii) auxiliary information, such as the parameters at initialization, through the state variable $s$. Naturally, not every regularizer takes all of these elements into account. The regularizer is optimized alongside the base objective, $J_{\tau(t)}(\theta)$. We write the composite objective as $J^{\lambda}_{\tau(t)}(\theta)=J_{\tau(t)}(\theta)+\lambda\mathcal{R}_{\tau(t)}(\theta,s)$, where $\lambda$ is a tunable hyperparameter governing the regularization strength.
One commonly used regularizer in both continual and non-continual learning is L2 regularization, $\mathcal{R}_{\tau(t)}(\theta,\varnothing)=\|\theta\|^{2}$, which regularizes the parameters towards zero. A recent alternative, L2 regularization towards the initialization, $\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\|\theta-\theta^{(0)}\|^{2}$, was proposed to deal with sensitivity to parameters near zero [17]. Regularizing towards the particular parameters present at initialization allows the neural network to regenerate parameters, providing a soft reset to parameters if the gradient on the base objective is zero (see Appendix A.3 for an example and Appendix A.5 for more details). One potential problem with L2 regularization towards the initialization is that it may prevent the parameters from deviating from the particular sampled value from the initialization distribution.
4.1 Wasserstein Regularization Towards Initialization
A first attempt to preserve the properties of the initialization is to regularize the parameter distribution rather than their specific values. Thus, we first propose a regularization that is simple and penalizes the parameters if they are far from the initialization distribution. At initialization, the parameters at layer $l$ are sampled i.i.d. $\theta_{i,j}\sim p^{(l,0)}(\theta)$ according to some pre-determined distribution $p^{(l,0)}$, such as the Glorot initialization [18] The distribution of parameters is only known at the initialization (at iteration 0); for any particular layer, the distribution $p^{(l,t)}$ is no longer has a known closed form at iteration $t$ during training. However, it is still possible to regularize the empirical distribution towards the initialization distribution by using the empirical Wasserstein-2 distance, defined as the difference in the order statistics [39]. We denote the flattened parameter matrix for layer $l$ at iteration $t$ by $\bar{\theta}^{(l,t)}$, and its order statistics by $\theta_{(i)}^{(l,t)}$ for $i=1,\ldots,d_{l}$, representing the $i$-th smallest parameter at iteration $t$ for layer $l$ (here $d_{l}$ denotes the number of parameters at layer $l$). The Wasserstein regularizer uses the empirical Wasserstein-2 distance for each layer of the neural network:
$\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\sum_{l=1}^{L}\sum_{i=1}^{d_{l}}\left(\bar{\theta}_{(i)}^{(l,t)}-\bar{\theta}_{(i)}^{(l,0)}\right)^{2}.$
In the above equation, we are taking the L2 difference between the order statistics of each layer’s parameters at initialization and at iteration $t$ during training. One of the benefits of using this distributional distance is that the parameters have more leeway to move away from the initialization compared to the L2 regularizer (see Appendix A.6 for details).
4.2 Spectral Regularization for Continual Learning
@@
In Section 3, we argued that the optimisation dynamics of continual learning may impede the continued trainability of a neural network. In particular, we showed that a growing parameter norm results in a growing spectral norm and condition number which can harm trainability by reducing effective gradient diversity. Now we investigate regularization as a means of controlling the spectral norm to maintain trainability of the neural network.
@@
The parameter matrix $\theta_{l}$ in layer $l$ is a composition of the weight matrix $\mathbf{W}_{l}$ and the bias $\mathbf{b}_{l}$, giving the augmented parameter matrix $\theta_{l}=[\mathbf{W}_{l}|\mathbf{b}_{l}]$ (with this the notation the equation defining the of the neural network $f_{\theta}$ can be rewritten as $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})$). As discussed in Section 3, our goal is to control the maximum singular value of $\theta_{l}$. Using that $\sigma_{max}(\theta_{l})=\sigma_{max}([\mathbf{W}_{l}|\mathbf{b}_{l}])\leq\sigma_{max}(\mathbf{W}_{l})+\sigma_{max}(\mathbf{b}_{l})$, we achieve this by regularizing the spectral norm of each parameter in the layer separately: the spectral norm of the multiplicative weight parameter $\mathbf{W}_{l}$ is regularized towards 1 and the spectral norm of the additive bias parameter $\mathbf{b}_{l}$ is regularized towards 0:
@@
$\mathcal{R}_{\tau(t)}(\theta_{l})=\sum_{l\in\text{layers}}\left((\sigma_{max}(\mathbf{W}_{l})^{k}-1)^{2}+\left(\sigma_{max}(\mathbf{b}_{l})^{k}\right)^{2}\right)=\sum_{l\in\text{layers}}\left((\sigma_{max}(\mathbf{W}_{l})^{k}-1)^{2}+\|\mathbf{b}_{l}\|_{2}^{2k}\right).$
@@
We add the exponent $k$ to the spectral norm to penalize large spectral norms deviating from one. In our experiments, we set $k=2$ and provide an ablation study in Appendix C.2. The maximum singular value can be computed efficiently using power iteration [40, 41]. Similarly to previous work using spectral regularization for generalization in supervised learning, we find that a single iteration is sufficient for effective regularization [42].
Spectral regularization explicitly regularizes the spectral norm (the maximum singular value) of each layer and addresses the problem of reduced effective gradient diversity described in Section 3.3. We only minimize the maximum singular value. We found that the smallest singular value remains relatively constant across all the architectures that we consider (see Appendix C.5). For implementation details for other layers, such as normalization layers and convolutional layers, see Appendix A.9.
@@
In addition to the exponent $k$, our approach to spectral regularization is different in two important ways compared to spectral-norm regularization [42]. First, we regularize the spectral norm of every parameter, including bias terms and normalization parameters (see Appendix A.9). This is required because every parameter experiences norm growth. Second, we regularize the spectral norm of the multiplicative parameters to one, rather than zero. Aggressive regularization of the spectral norm towards zero can lead to collapse issues similar to L2 regularization towards zero. These two differences are crucial to improving trainability in continual learning. Lastly, spectral regularization is preferable over spectral normalization for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 in [43], and Appendix A.8), which can be problematic in continual learning due to the changing data distribution, and (ii) other forms of normalization are already used to train deep neural networks, such as LayerNorm [44], and it has been shown that spectral normalization does not improve trainability [5]. See Appendix A.7 for more details.
@@ 5 Experiments
Our experiments are designed to investigate the effect of both Wasserstein and spectral regularization on trainability in continual supervised learning. They cover several different datasets, types of nonstationarity, and architectures. We compare our proposed regularizers against baselines that have shown to improve trainability in previous work, which we detail below. Overall, our experiments demonstrate that spectral regularization, and to a lesser extent Wasserstein regularization (i) consistently mitigates loss of trainability on a wide variety of continual supervised learning problems, including training large neural networks for thousands of epochs across hundreds of tasks; (ii) is highly robust to the number of training epochs per task, type of non-stationarity, and choice of regularization strength; and (iii) achieves better generalization performance over the course of continual learning.
@@
Our experiments use the following datasets for continual supervised learning: MNIST [45], Fashion MNIST [46], EMNIST [47], SVHN2 [26], CIFAR10, and CIFAR100 [27]. In addition to the dataset, we consider three different types of non-stationarity: (i) random label assignments, (ii) pixel permutation, and (iii) label flipping. Random label assignments are commonly used to evaluate trainability [5] due to the large distribution shift between tasks in memorizing completely new and random labels. Pixel permutations, on the other hand, require only learning the permutation mask applied to the image; it induces loss of trainability more slowly but is useful for evaluating generalization [17]. Lastly, label flipping is a re-assignment of all the observations from one label to another label. Unlike the other two non-stationarities, a label flip distribution shift does not require learning a new representation. This is because only the output layer of a neural network needs to be permuted to learn the label re-assignment, but gradients can still unnecessarily change the representation leading to loss of trainability [12]. For the experiments on MNIST, Fashion MNIST and EMNIST, we use feed-forward neural networks. On SVHN2, CIFAR10, and CIFAR100 we use both a ResNet-18 [25] and Vision Transformer [24].
@@
In our main results, we compare Wasserstein regularization and spectral regularization against L2 regularization towards zero, L2 regularization towards the initialization [17], recycling dormant neurons [23], shrink and perturb [15] and no regularization. In Appendix C.1, we provide additional results on other baselines: AdamW, which also regularizes towards zero and concatenated ReLU [4, 48]. Because we are interested in trainability in the task-agnostic setting, we do not compare with regularizers for catastrophic forgetting as they rely on privileged information regarding the task [49, 50]. Our experiments largely use Layer Norm [44] or Batch Norm in the case of ResNet-18 [51], which has been shown to be a mitigator for loss of trainability [37]. We also provide an ablation study showing that Layer Norm is synergistic with the regularizers we propose.
@@ 5.1 Comparative Evaluation
Figure 1: Comparison of task accuracy with random label assignment. Spectral regularization is particularly effective at mitigating loss of trainability. Without regularization, trainability is always reduced.
In Figure 1, we plot the results of training a ResNet-18 on different image classification datasets with random label assignment as the type of non-stationarity for each task. Across all problems and methods considered, spectral regularization was best at improving and sustaining its trainability, with accuracy increasing as new tasks were encountered. In contrast, despite improved trainability over the first few tasks, the unregularized baseline eventually experienced loss of trainability, leading to lower accuracy. Recycling dormant neurons, which reinitializes weights based on dormant neurons, experienced loss of trainability on all problems. One potential reason for this is because this approach only resets the weights of the convolutional layers and the final fully-connected layer [23], whereas our approach regularizes all parameters, including bias terms and batch normalization parameters. In Appendix C.1, we show that weight reinitialization can be effective on smaller problems without normalization layers. Shrink and perturb is similar to L2 regularization towards zero, but injects noise and is not as effective. While the perturbations were previously demonstrated to improve generalization when warm-starting a neural network on a new problem, we see that they are not effective when training on a longer sequence of problems. Wasserstein regularization is often on par with L2 regularization towards the initialization, and surpasses it on CIFAR10. These results demonstrate that, in contrast to other approaches, spectral regularization consistently maintains trainability across problems.
@@ 5.2 Sensitivity Analysis
In this section, we investigate the robustness of spectral regularization when varying (i) the number of training epochs per task, (ii) the type of non-stationarity applied, or, (iii) the regularization strength.
Varying the Number of Epochs Per Task
Given that the continual learning problem outlined in Section 2 depends on the number of iterations, it may be the case that a higher number of epochs per task can mitigate loss of trainability. In Figure 2, we found that the number of epochs only delays loss of trainability (see Appendix C.3 for results on Fashion MNIST). Even when the number of epochs per task is high enough to reach 100% accuracy on the first task, loss of trainability eventually occurred without regularization. In contrast, when using spectral regularization, loss of trainabililty was consistently mitigated. We additionally found that spectral regularization is particularly effective at maintaining trainability when the number of epochs per tasks is low. The evidence for this is most striking without Layer Norm, in the top row, where spectral regularization was the only method capable of maintaining its trainability. Even though Layer Norm controls the spectral norm through the activations [52], adding Layer Norm alone was not enough to mitigate loss of trainability. However, Layer Norm is synergistic with various forms of regularization, with accuracy being improved regardless of the number of epochs per task.
Figure 3: Sensitivity to type of non-stationarity on EMNIST. Both Wasserstein and spectral regularization are able to consistently maintain high trainability across random label assignment, pixel permutations, and label flipping. When applying random label assignment, L2 (init) and L2 (zero) are unable to attain high trainability as compared to Wasserstein and spectral regularization.
Varying Type of Non-stationarity
Next, we explore whether our proposed regularizers are effective across different types of non-stationarity. To this end, we use EMNIST because the number of classes is large enough that label flipping induces loss of trainability [12]. In Figure 3, we found that both Wasserstein and spectral regularization best maintained trainability across different non-stationarities.
Figure 4: Sensitivity to regularization strength on MNIST with random label assignment. Spectral regularization is insensitive to regularization strength while remaining performant at few epochs.
Varying Regularization Strength
@@
The effectiveness of a regularizer depends on the regularization strength. Too much regularization can slow down training, leading to suboptimal performance on any given task. However, too little regularization can lead to loss of trainability and suboptimal performance on future tasks. In a continual learning problem, hyperparameter tuning is particularly costly (or infeasible), especially if the distribution over tasks is not known in advance [53]. Thus, we should favour regularizers that are less sensitive to hyperparameters. Our results in Figure 4 show that spectral regularization dominates other regularizers for any choice of regularization strength, is much highly insensitive to the choice of hyperparameter compared to other regularizers.
Figure 5: Effects of regularization on distance from initialization, average condition, average weight rank, and gradient diversity when training ResNet-18 on SVHN2 with random label assignment. Despite being far from initialization, spectral regularization has a low average condition number, a high effective rank across its parameters, and a high effective gradient diversity.
5.3 Looking Inside the Network
We now explore structural properties of the neural network, and how these properties are effected by our proposed regularizers. In Figure 5, we further confirm that spectral regularization is the least restrictive approach, allowing deviations far from the initialization. Spectral regularization also controlled the spectral norm of the network, leading to a better average condition number (the ratio between the maximum and the minimum singular values), a higher effective rank of its parameters, and a higher gradient diversity. See Appendix C.5 for similar results on the other datasets considered.
Figure 6: Comparison of generalization performance with random label assignment. Spectral regularization improves generalization, using either a ResNet-18 (Top) or a Vision Transformer (bottom).
5.4 From Trainability to Generalization
Having demonstrated the main goal of this paper, namely the effectiveness of Wasserstein and spectral regularization on trainability, we now investigate generalization. We used a pixel permutation non-stationarity in Figure 6, and found that spectral regularization is better able to generalize compared to other regularizers, using either a ResNet-18 or Vision Transfomer. Interestingly, we also find that the Vision Transformer, particularly with spectral regularization, provides better generalization performance than using a ResNet. This demonstrates that regularizers which more directly address the issues underlying trainability also lead to better generalization.

### v2

@@ 4 Spectral Regularization for Continual Learning
If important properties of the initialization are lost during the course of learning, it is natural to regularize the neural network toward the initialization. This is the motivation for regularization (Kumar et al., 2023) and weight reinitialization (Sokar et al., 2023, Dohare et al., 2024) as loss of plasticity mitigators. However, our motivation is to more directly target key properties of initialization, using the insights from Section 3.
@@
We denote a regularizer by $\mathcal{R}_{\tau(t)}(\theta,s)$, which is a function of (i) the parameters of the neural network, $\theta$, (ii) the data, through the current task $\tau(t)$, and of (iii) auxiliary information, such as the parameters at initialization, through the state variable $s$. Naturally, not every regularizer takes all of these elements into account. The regularizer is optimized alongside the base objective, $J_{\tau(t)}(\theta)$. We write the composite objective as $J^{\lambda}_{\tau(t)}(\theta)=J_{\tau(t)}(\theta)+\lambda\mathcal{R}_{\tau(t)}(\theta,s)$, where $\lambda$ is a tunable hyperparameter governing the regularization strength.
@@
In Section 3, we argued that a growing spectral norm and condition number can harm trainability by reducing effective gradient diversity and that this growth occurs at a rate of $\sqrt{t}$. Now we investigate regularization as a means of controlling the spectral norm to maintain the trainability of the neural network.
One commonly used regularizer in both continual and non-continual learning is L2 regularization, $\mathcal{R}_{\tau(t)}(\theta,\varnothing)=\|\theta\|^{2}$, which regularizes the parameters towards zero. A recent alternative, L2 regularization towards the initialization, $\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\|\theta-\theta^{(0)}\|^{2}$, was proposed to deal with sensitivity to parameters near zero (Kumar et al., 2023). Regularizing towards the particular parameters present at initialization allows the neural network to regenerate parameters, providing a soft reset to parameters if the gradient on the base objective is zero (see Appendix A.2 for an example and Appendix A.4 for more details). One potential problem with L2 regularization towards the initialization is that it may prevent the parameters from deviating from the particular sampled value from the initialization distribution.
@@
Our proposed spectral regularizer explicitly regularizes each layer’s spectral norm (the maximum singular value) and addresses the reduced effective gradient diversity described in Section 3.3. We only minimize the maximum singular value because the smallest singular value remains relatively constant across all the architectures we consider (see Appendix C.6). The parameter matrix $\theta_{l}$ in layer $l$ is a concatenation of the weight matrix $\mathbf{W}_{l}$ and the bias $\mathbf{b}_{l}$, giving the augmented parameter matrix $\theta_{l}=[\mathbf{W}_{l}|\mathbf{b}_{l}]$ and augmented input $\tilde{h}_{l}=[h_{l},\mathbf{1}]$. With this notation, the equation defining the neural network $f_{\theta}$ can be rewritten as $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})=\phi(\theta_{l}\tilde{h}_{l})$. As discussed in Section 3, our goal is to control the largest singular value of $\theta_{l}$. Using the fact that the concatenated parameters can be upper-bounded by the sum, $\sigma_{1}(\theta_{l})=\sigma_{1}([\mathbf{W}_{l}|\mathbf{b}_{l}])\leq\sigma_{1}(\mathbf{W}_{l})+\sigma_{1}(\mathbf{b}_{l})$, we achieve this by regularizing the spectral norm of each parameter in the layer separately: the spectral norm of the multiplicative weight parameter $\mathbf{W}_{l}$ is regularized towards one and the spectral norm of the additive bias parameter $\mathbf{b}_{l}$ is regularized towards zero:
@@
$\mathcal{R}_{\tau(t)}(\theta_{l})=\sum_{l\in\text{layers}}\left((\sigma_{1}(\mathbf{W}_{l})^{k}-1)^{2}+\left(\sigma_{1}(\mathbf{b}_{l})^{k}\right)^{2}\right)=\sum_{l\in\text{layers}}\left((\sigma_{1}(\mathbf{W}_{l})^{k}-1)^{2}+\|\mathbf{b}_{l}\|_{2}^{2k}\right).$
@@
We add the exponent $k$ to the spectral norm to penalize large spectral norms deviating from one. We set $k=2$ in our experiments and provide an ablation study in Appendix C.1. For implementation details for other layers, such as normalization and convolutional layers, see Appendix A.7. The largest singular value can be computed efficiently using power iteration (Golub and Van der Vorst, 2000, Householder, 2013). Similar to previous work using spectral regularization for generalization in supervised learning, we find that a single iteration is sufficient for effective regularization (Yoshida and Miyato, 2017).
@@
In addition to the exponent $k$, our approach to spectral regularization is different in two important ways compared to previous work, such as spectral-norm regularization (Yoshida and Miyato, 2017). First, we regularize the spectral norm of every parameter, including bias terms and normalization parameters (see Appendix A.7). This is required because every parameter experiences norm growth. Second, we regularize the spectral norm of the multiplicative parameters to one rather than zero. Aggressive regularization of the spectral norm towards zero can lead to collapse issues similar to L2 regularization towards zero. These two differences are crucial to improving trainability in continual learning. Lastly, spectral regularization is preferable over spectral normalization, which explicitly normalizes the weights in the forward pass to be exactly one, for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 by Miyato et al. (2018), and Appendix A.6), which can be problematic in continual learning due to the changing data distribution, and, (ii) other forms of normalization are already used to train deep neural networks, such as LayerNorm (Ba et al., 2016), and it has been shown that spectral normalization does not improve trainability (Lyle et al., 2023). See Appendix A.5 for more details.
@@ 5 Experiments
The goal of our experiments is to investigate the effect of spectral regularization on trainability in continual supervised learning, as well as reinforcement learning. We cover several different datasets, types of nonstationarity, and architectures. We compare our proposed regularizers against baselines that have been shown to improve trainability in previous work, which we detail below. Overall, our experiments demonstrate that spectral regularization (i) consistently mitigates loss of trainability on a wide variety of continual supervised learning problems, including training large neural networks for thousands of epochs across a hundred tasks, (ii) is highly robust to the regularization strength, type of non-stationarity and the number of training epochs per task, (iii) achieves better generalization performance over the course of continual learning, and (iv) is generally applicable, which we demonstrate by applying spectral regularization to reinforcement learning with continuous actions.
@@
Our main results uses all commonly used image classification datasets for continual supervised learning: tiny-ImageNet (Le and Yang, 2015), CIFAR10, and CIFAR100 (Krizhevsky, 2009). Experiments in the appendix also use smaller-scale datasets, like MNIST (LeCun et al., 1998), Fashion MNIST (Xiao et al., 2017), EMNIST (Cohen et al., 2017), and SVHN2 (Netzer et al., 2011). In addition to the dataset, we consider different types of non-stationarity: (i) random label assignments, (ii) pixel permutation, (iii) label flipping, and (iv) class-incremental learning. Random label assignments are commonly used to evaluate trainability (Lyle et al., 2023) due to the large distribution shift between tasks in memorizing completely new and random labels. Pixel permutations, on the other hand, require only learning the permutation mask applied to the image; it induces loss of trainability more slowly but is useful for evaluating generalization (Kumar et al., 2023). Label flipping is a re-assignment of all the observations from one label to another label. Unlike the other two non-stationarities, a label flip distribution shift does not require learning a new representation. This is because only the output layer of a neural network needs to be permuted to learn the label re-assignment, but gradients can still unnecessarily change the representation leading to loss of trainability (Elsayed and Mahmood, 2024). We also consider the class-incremental setting in which the network is trained on a growing subset of the classes from a dataset, starting with only five classes on the first task and introducing five new classes on new tasks. We use both a ResNet-18 (He et al., 2016) and Vision Transformer (Dosovitskiy et al., 2021).
@@
In our main results, we compare spectral regularization against L2 regularization towards zero, shrink and perturb (Ash and Adams, 2020), L2 regularization towards the initialization (Kumar et al., 2023), recycling dormant neurons (ReDO, Sokar et al., 2023), concatenated ReLU (Abbas et al., 2023, Shang et al., 2016), and Wasserstein regularization (Lewandowski et al., 2023). Several regularizers in the continual learning without forgetting literature rely on privileged task information, which is not applicable to the task-agnostic setting that we consider. We use the streaming conversion (Elsayed and Mahmood, 2024) to transform elastic weight consolidation (Kirkpatrick et al., 2017, Zenke et al., 2017), so that it no longer requires task boundary information, and include it as a baseline. Additional experiment details can be found in Appendix B.
@@ 5.1 Comparative Evaluation
Figure 1: Generalization across different types of non-stationarity on tiny-ImageNet using a ResNet (top) or a Vision Transformer (bottom). Compared to the baselines, spectral regularization is consistently among the best-performing methods across class incremental, label flip, and pixel permutation non-stationarities. Note that the Vision Transformer often achieves better generalization performance than the ResNet architecture.
In Figure 1, we plot the results of training a ResNet-18 and a Vision Transformer on tiny-ImageNet with different non-stationarities. Across networks, non-stationarities, and methods considered, we see that spectral regularization is among the methods best capable of sustaining plasticity. The advantage of our approach, spectral regularization, is particularly high in class-incremental learning, but the performance of spectral regularization with a Vision Transformer was also the best on label flipping and pixel permutation. In contrast, the performance of other baselines was highly variable with respect to the specific type of non-stationarity. For example, shrink and perturb is not robust to the settings considered. Sometimes, it is among the best-performing or worst-performing methods. Similar results on other datasets and the pixel permutation non-stationarity can be seen in Figure 2.
Figure 2: Continual learning with pixel permutation tasks on SVHN2, CIFAR10, CIFAR100 using a ResNet-18 (top) or a Vision Transformer (bottom). Across different datasets, spectral regularization is effective at maintaining test accuracy on new tasks. Without any mitigators, both ResNet-18 and Vision Transformer have diminishing test accuracy, suggesting loss of plasticity.
5.2 Looking Inside the Network
Figure 3: Trainability and neural network properties across ImageNet, CIFAR10, and CIFAR100. Baselines that suffer from a loss of trainability (top) also have an increasing average spectral norm (middle-top), and a decrease in their average representation change (middle-bottom).
We now explore how the structural properties of a neural network evolve over the course of continual learning, how these properties are affected by spectral regularization, and the different baselines considered. For this, we consider the continual learning problem in which a ResNet-18 must memorize a set of random labels which changes from task to task using the tiny-ImageNet, CIFAR10, and CIFAR100 datasets. We consider the average representation change to measure the distance between the neural network’s hidden activations from the beginning of one task, to the beginning of the next task. The average representation change is a proxy for plasticity, allowing us to see how much the behavior of the neural network has changed. In Figure 3, we see that the unregularized ReLU networks suffers from loss of trainability in all problems considered, and this coincides with an increasing spectral norm (middle-top) and a decrease in the average representation change (bottom). Although only spectral regularization directly regularizes the spectral norm of the network, other regularizers do so indirectly by controlling other norms, like L2. However, these other regularizers also regularize other parameters, preventing them from deviating from initialization and potentially leading to suboptimal use of capacity, which can be observed in the bottom row in Figure 3. In Appendix C.3, we show that the Vision Transformer was not able to memorize random labels, suggesting that its capabilities for generalization are offset by its relatively lower trainabilty.
@@ 5.3 Sensitivity Analysis
Figure 4: Sensitivity analysis on regularization strength. Compared to other regularizers, spectral regularization is insensitive to regularization strength while sustaining higher trainability for any given regularization strength.
@@
The effectiveness of a regularizer depends on the regularization strength. Too much regularization can slow down training, leading to suboptimal performance on any given task. However, too little regularization can lead to loss of trainability and suboptimal performance on future tasks. In a continual learning problem, hyperparameter tuning is particularly costly (or infeasible), especially if the distribution over tasks is not known in advance (Mesbahi et al., 2024). Thus, we should favour regularizers that are less sensitive to hyperparameters. Our results in Figure 10 show that the sensitivity of spectral regularization to its hyperparameter is much lower than other regularizers. In Appendix C.4, we also present additional results showing the robustness of spectral regularization when varying (i) the regularization strength using other network architectures, (ii) the type of non-stationarity, and (iii) the number of training epochs per task.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 5: Spectral regularization enhances plasticity in reinforcement learning in the DMC suite. Spectral regularization is competitive with the network reset + layernorm (Reset) even when the replay buffer is unbounded (Top Left). When the replay buffer size is bounded to 250k steps, spectral regularization improves over the Reset baseline (Bottom Left). In both cases, Spectral regularization significantly outperforms layernorm (Baseline), Compared to the hyperparameter governing the reset frequency, spectral regularization is less sensitive to its regularization strength. Spectral regularization also prevents both parameter and gradients from exploding, and reduces value overestimation (Right).
5.4 From Supervised Learning to Reinforcement Learning
In addition to supervised learning, we evaluate spectral regularization in reinforcement learning. Specifically, we investigate the control tasks from the DMC benchmark (Tassa et al., 2020), with soft actor-critic  (SAC, Haarnoja et al., 2018), in the high replay-ratio (RR) regime (D’Oro et al., 2022) with 16 gradient updates for every environment step. The high replay-ratio regime leads to primacy bias (Nikishin et al., 2022), a phenomenon related to loss of plasticity. Recent work has demonstrated that only full network resets with layer normalization (Ba et al., 2016) serve as an effective mitigation strategy in this setting (Ball et al., 2023, Nauman et al., 2024).
As shown in Figure 5, spectral regularization yields substantially higher returns compared to the baseline that uses layer normalization with either an unbounded replay buffer size or a limited replay buffer size. In aggregate, spectral regularization is competitive with the final performance of the strongest baseline, full network resets in combination with layer normalization. However, in five out of seven environments, spectral regularization substantially improves over the reset baseline in terms of sample efficiency (see Appendix C.8). Spectral regularization is also less sensitive to its regularization strength than resets are for their reset frequency (Figure 5, middle-top and middle-bottom). Moreover, we used a single regularization strength for every network. Better performance may be achieved by individually tuning each regularization strength for the value and policy networks. In addition, a combination of spectral normalization and layer norm aggressively reduces Q-value overestimation (Figure 5, top-left), which is one of the well-established proxies for RL training destabilization (Hasselt, 2010, Nauman et al., 2024). Lastly, spectral normalization prevents exploding gradients and keeps weights’ norms small during training, which is a key component for learning continually and preserving the plasticity of neural networks (Lyle et al., 2023; 2022).